# Learning Neural Differential Algebraic Equations via Operator Splitting

### v2

@@ Abstract
Differential-Algebraic Equations (DAEs) describe the temporal evolution of systems that obey both differential and algebraic constraints. Of particular interest are systems that contain implicit relationships between their components, such as conservation relationships. Here, we present an Operator Splitting (OS) numerical integration scheme for learning unknown components of Differential-Algebraic Equations from time-series data. This methodology is built upon the concept of the Universal Differential Equation; that is, a model constructed as a system of Neural Ordinary Differential Equations informed by theory from particular science domains. In this work, we show that the proposed OS-based time-stepping scheme is suitable for relevant system-theoretic data-driven modeling tasks. Presented examples include (i) the inverse problem of tank-manifold dynamics and (ii) discrepancy modeling of a network of pumps, tanks, and pipes. Our experiments demonstrate the proposed method’s robustness to noise and extrapolation ability to (i) learn the behaviors of the system components and their interaction physics and (ii) disambiguate between data trends and mechanistic relationships contained in the system.
@@ I INTRODUCTION
Modeling and simulation of systems via Differential-Algebraic Equations (DAEs) can be difficult, both in terms of formulation as well as numerical implementation. Common issues include (i) over- or under-constraining states and (ii) incorrect model specification, either of which can lead to a prohibitively challenging modeling task [1]. While many such algebraic relationships may be known (e.g., mass is conserved through a pipe), incorporating these algebraic relationships into well-established data-driven modeling paradigms is non-trivial. Here, measurement noise, partial observations, and missing or incorrect physics confound the ability to effectively build data-driven models that obey algebraic constraints. In practice, simplified models can be adopted that capture improperly specified or missing physics at a given fidelity. A common approach is to employ a closure model [2], i.e., a surrogate for the full-fidelity physics inserted into a lower-fidelity model. Closure models have enabled major advancements in scientific computing but still traditionally rely upon rigorous reduction from first principles physics to derive required simplified relationships [3].
@@ I-A Contributions
  • Taking inspiration from the well-established field of operator splitting, we extend neural timesteppers to appropriately integrate DAEs.
  • We accomplish this via an algebra surrogate layer that modifies algebraic states of a system, while an ODE Solver integrates differential states.
  • We demonstrate the utility of the method in data-driven parameter estimation problems such as canonical closure and inverse modeling of constrained networked system dynamics.
@@ I-B1 Machine Learning for Dynamical Systems
Methods for learning dynamics from data exist on a continuum from ‘black-box’ models to ’white-box’ models [8]. The ‘black-box’ modeling task is performed without regard to the underlying system or domain knowledge and includes Recurrent Neural Networks (RNNs) and their variants (e.g., LSTMs) [9], neural state space models [10, 11], neural timesteppsers [4], sparse symbolic regression [12], and Neural ODEs (NODEs) [7, 13, 14]. Often, such models extract salient features in the training data to infer future states. However, satisfying algebraic constraints in these data-driven models is not trivial. ‘White-box’ models use domain knowledge to constrain models to match expected dynamics, often leading to parameter-tuning problems [15]. In ‘gray-box’ models, recently referred to as physics-informed machine learning [16], prior knowledge is incorporated in a manner to constrain a ‘black-box model’. Neural ODEs constructed with domain knowledge, termed Universal Differential Equations [17, 18], are an attractive modeling platform for these problems because of their flexibility in encoding structural priors.
@@ I-B2 Data-driven Modeling of DAEs
While the specific treatment of each of these domain-focused problems varies, each leverages a form of a multi-step integration technique that separates the evolution of differential and algebraic states. We recognize that the commonality of these works holds the potential for a generalization of standard NODE integration. Specifically, we recast these ideas under operator splitting methods, where integration is split into simpler, sequential sub-tasks. In effect, we propose a new class of neural timesteppers suitable for the data-driven modeling of DAEs with partially unknown dynamics.
@@ II-B Neural Differential Equations
where the superscript contained in parentheses denotes indexing by time. This integration is performed with a standard numerical ODE solver, which we denote as:
$x^{(t+\Delta t)}=\operatorname{ODESolve}\left(f,x^{(t)};\theta\right),$ (4)
where the limits of temporal integration are understood to be $[t,t+\Delta t]$. In general, the NODE concept extends $f$ to any differentiable mapping by implementing Eq. (4) in a differentiable manner [17]. This is achieved through either (i) backpropagating residuals through the elementary functions of the explicit time-stepping integrator (reverse-mode automatic differentiation) [26, 4], (ii) an adjoint sensitivity definition and solution [7], or (iii) forward-mode automatic differentiation [27].
@@ III PROBLEM FORMULATION
A fundamental assumption of a NODE model (2) is that a direct relationship exists between each system state and its temporal derivative. This is a reasonable assumption for many cases, but poses challenges for partially-observed systems [28] and application to DAEs. Index reduction is a reliable technique for recasting a DAE into a set of ODEs for standard numerical integration [29], but this is incompatible with the problem when the DAE governing equations are unknown or incomplete, leading to an unknown system of ODEs and required state augmentations.
@@
$\underset{\Theta}{\text{min}}$ $\int_{t_{0}}^{t_{N}}\Big{(}||x-\hat{x}||_{2}^{2}+||y-\hat{y}||_{2}^{2}\Big{)}$ (5a)
s.t $\frac{dx}{dt}=f\left(x,y,u;\Theta\right),$ (5b)
$0=g\left(x,y,u;\Theta\right),$ (5c)
@@
Where $\hat{x},\hat{y}$ represent reference state variables generated by the unknown DAE system (1). The objective of the problem (5a) is to train the parameters $\Theta$ such that the reference state trajectories in the time interval $[{t_{0}},{t_{N}}]$ are closely approximated by the NDAE model given via Eq. (5b) and Eq. (5d), respectively.
@@
Under the above assumptions, there exists a unique, continuously differentiable solution $(x^{(t)},y^{(t)})$ defined on an interval $[t_{0},t_{0}+\Delta t)$, for some $\Delta t>0$, such that the DAE system is satisfied and $(x^{(t_{0})},y^{(t_{0})})=(x_{0},y_{0})$. Furthermore, the algebraic constraint can be locally solved for $y$ as a function of $x$ and $u$, thereby reducing DAEs into a system of ODEs $\frac{\partial x}{\partial t}=f(x,h(x,u),u)$ to which the classical Picard-Lindelöf Theorem III.2 applies.
Theorem III.2 (Classical Picard-Lindelöf for ODEs [31])
Let $f:D\subseteq\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{u}}\to\mathbb{R}^{n_{x}}$ be a function that is continuous in both arguments and Lipschitz continuous with respect to the state $x$, uniformly in $t$, on an open set $D$ containing the point $(t_{0},x_{0})$. Then there exists an interval $I=[t_{0}-\epsilon,t_{0}+\epsilon]\subset\mathbb{R}$, for some $\epsilon>0$, and a unique continuously differentiable function $:I\to\mathbb{R}^{n_{x}}$ such that
$\frac{\partial x(t)}{\partial t}=f(x(t),u(t)),\quad x(t_{0})=y_{0},$
for all $t\in I$. In other words, the ODE initial value problem has a unique local solution.
@@ IV METHOD
IV-A Operator Splitting Methods
Our methodology is inspired by operator splitting methods (OS) for evolving complex multiphysics problems. In OS, also called fractional step methods, the evolution of a system is split into simpler, independent sub-tasks. The underlying ODE is then solved explicitly as the sequential solution of these sub-tasks, albeit at a decreased level of accuracy (i.e., splitting error [32]). There are several well-established techniques for decomposing the solution of an ODE into sub-tasks. The simplest scheme, Lie-Trotter operator splitting [5], assumes a differential operator can be decomposed into two sub-tasks. For the ODE of Eq. (2), assume $f$ can be decomposed into operators $A$ and $B$ such that:
$\frac{dx}{dt}=A(x)+B(x),$ (6)
for the interval $t\in[t,t+\Delta t]$. The Lie-Trotter splitting scheme defines the sub-tasks as:
$\hat{x}^{(t+\Delta t)}$ $=\operatorname{ODESolve}\left(A(x),x^{(t)}\right),$ (7)
$x^{(t+\Delta t)}$ $=\operatorname{ODESolve}\left(B(x),\hat{x}^{(t+\Delta t)}\right).$ (8)
The first sub-task solves the ODE defined by the operator $A(x)$ on the time interval $[t,t+\Delta t]$. The second sub-task solves the ODE defined by the operator $B(x)$ on the same time interval, but with the initial conditions defined by the solution of the first sub-task. Many variations of this scheme exist for different applications and orders of accuracy. For example, Strang splitting [33] is a second-order scheme that uses half-steps (i.e., steps of $\frac{\Delta t}{2}$) in the above scheme.
IV-B Operator Splitting for Differential Algebraic Equations
Now consider the DAE system of Eq. (1). Similar to the approach of a Lie-Trotter splitting scheme, we split the evolution of a system’s differential and algebraic states into sequential sub-tasks. The first sub-task updates the algebraic states using a trainable deep neural network (DNN). The second sub-task takes as input the updated algebraic states and performs a single-step integration. These two sequential sub-tasks define one complete forward pass; all system states are updated exactly once. More formally, consider a differentiable ODE layer, denoted as $\operatorname{ODESolve}$. Using the operator splitting method as our solution technique for Eq. (1), we formulate a time stepper for Neural DAE as:
$y^{(t+\Delta t)}$ $=h\left(x^{(t)},y^{(t)},u^{(t)};\theta_{h}\right)$ (9a)
@@ IV-B Operator Splitting for Differential Algebraic Equations
where $h:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{y}}\times\mathbb{R}^{n_{u}}\rightarrow\mathbb{R}^{n_{y}}$ pairs with $g$ as an explicit algebra solver or trainable neural surrogate thereof and the temporal integration bounds are assumed to be $[t,t+\Delta t]$. We denote this complete set of operations as $\operatorname{DAESolve}$, with one complete time step captured as:
@@
where $\Theta=\{\theta_{h},\theta_{f}\}$ is the set of all model parameters.
Note that in this formulation, the updates to the algebraic and differential states are sequential; that is, the $\operatorname{ODESolve}$ step is dependent upon the update to the algebraic states. The sub-tasks do not commute. This is an essential assumption that enables explicit updates for both algebraic states and differential states. Figure 1 depicts the handling of differential and algebraic states through the proposed OS-based NDAE integration scheme. The module generates an internal state $y^{(t+\Delta t)}$ that is processed by the subsequent $\operatorname{ODESolve}$ as an exogenous input, similar to the notion of Neural Controlled Differential Equations [34].
@@
IV-C Loss Function and Optimization Algorithm
@@
We optimize model parameters $\Theta$ with an off-the-shelf stochastic gradient-based optimizer [35] to solve $\Theta^{*}=\underset{\Theta}{\text{argmin}}~{}\mathcal{L}(\Theta)$. For this work, we choose minimal architectures for ease of implementation and exposition, as outlined in Alg. 1. We use a first-order explicit timestepper for the ODE integration and use a DNN as a neural surrogate of an algebra solver. However, the proposed OS-based NDAE scheme is more generic than that and allows the combination of higher-order ODE solvers with iterative algebraic solvers such as Newton’s method.
@@ V-A1 Problem Description
A manifold is a physical realization of a conservation relationship that manifests as an algebraic constraint in a DAE. Figure 2 depicts a pair of tanks fed by a manifold. A pseudo-infinite reservoir supplies the pump. At the manifold, the inflow is exactly equal to the sum of the outflows; that is, $y_{\text{in}}=y_{1}+y_{2}$, where $y$ represents volumetric flow rates and the subscripts $1$ and $2$ denote the two exit streams of the manifold. In this example, the two tanks have different area-height profiles but have a common datum such that the column heights are identical. As a fluid is pumped into the two tanks, the flow rates in the two streams varies to ensure (i) the conservation relationship of the manifold and (ii) that each of the tanks have identical pressure head, i.e $x_{1}=x_{2}$, where $x$ is the height of the fluid in the respective tank. The full DAE that describes the evolution of this system is given as:
$\begin{split}\frac{dx_{1}}{dt}&=\frac{y_{1}}{\phi_{1}(x_{1})}\\ \frac{dx_{2}}{dt}&=\frac{y_{2}}{\phi_{2}(x_{2})}\end{split}$ (12)
$0=y_{\text{in}}-y_{1}-y_{2}\quad$ $0=x_{1}-x_{2},$
where $\phi_{1}$ and $\phi_{2}$ denote the area-height profiles for the two tanks. The solution of Eq. (12) can be obtained with many off-the-shelf standard scientific computing packages. For this study, the area-height profiles are set to $\phi_{1}(x)=3.0$ and $\phi_{2}(x)=\sqrt{x}+0.1$ respectively. The goal of the learning task is to recover $\phi_{2}(x)$ given observations of tank heights $x$ and pump volumetric flow rates $y$. Note that if the task were to recover both $\phi_{1}(x)$ and $\phi_{2}(x)$ simultaneously, the problem would be ill-posed in the sense that there does not exist a unique set of tank profiles that recovers the behavior seen in observations.
@@
Figure 2: Tank-manifold-pump schematic. A reservoir supplies fluid to a pump-manifold system. The tanks have different area-height profiles but a common datum. Conservation is enforced at the manifold.

### v3

@@ Abstract
Differential algebraic equations (DAEs) describe the temporal evolution of systems that obey both differential and algebraic constraints. Of particular interest are systems that contain implicit relationships between their components, such as conservation laws. Here, we present an Operator Splitting (OS) numerical integration scheme for learning unknown components of DAEs from time-series data. In this work, we show that the proposed OS-based time-stepping scheme is suitable for relevant system-theoretic data-driven modeling tasks. Presented examples include (i) the inverse problem of tank-manifold dynamics and (ii) discrepancy modeling of a network of pumps, tanks, and pipes. Our experiments demonstrate the proposed method’s robustness to noise and extrapolation ability to (i) learn the behaviors of the system components and their interaction physics and (ii) disambiguate between data trends and mechanistic relationships contained in the system.
@@ I INTRODUCTION
Modeling and simulation of systems via differential algebraic equations (DAEs) can be difficult, both in terms of formulation as well as numerical implementation. Common issues include (i) over- or under-constraining states and (ii) incorrect model specification, either of which can lead to a prohibitively challenging modeling task [1]. While many such algebraic relationships may be known (e.g., mass is conserved through a pipe), incorporating these algebraic relationships into well-established data-driven modeling paradigms is non-trivial. Here, measurement noise, partial observations, and missing or incorrect physics confound the ability to effectively build data-driven models that obey algebraic constraints. In practice, simplified models can be adopted that capture improperly specified or missing physics at a given fidelity. A common approach is to employ a closure model [2], i.e., a surrogate for the full-fidelity physics inserted into a lower-fidelity model. Closure models have enabled major advancements in scientific computing but still traditionally rely upon rigorous reduction from first principles physics to derive required simplified relationships [3].
@@ I-A Contributions
  • Taking inspiration from the field of operator splitting, we extend neural timesteppers to integrate DAEs. We achieve that via a neural surrogate for algebraic constraints alongside an ODE solver for differential states.
  • We demonstrate the method’s utility in data-driven parameter estimation tasks, including closure and inverse modeling of constrained network dynamics.
@@ I-B1 Machine Learning for Dynamical Systems
Methods for learning dynamics from data exist on a continuum from black-box models to white-box models [8]. The black-box modeling task is performed without regard to the underlying system or domain knowledge and includes Recurrent Neural Networks (RNNs) and their variants (e.g., LSTMs) [9], neural state space models [10, 11], neural timesteppsers [4], Neural ODEs (NODEs) [7, 12, 13], sparse symbolic regression [14], and Koopman operator [15]. Often, such models extract salient features in the training data to infer future states. However, satisfying algebraic constraints in these data-driven models is not trivial. White-box models use domain knowledge to constrain models to match expected dynamics, often leading to parameter-tuning problems [16]. In gray-box models, recently referred to as physics-informed machine learning [17], prior knowledge is incorporated in a manner to constrain a black-box model. Neural ODEs constructed with domain knowledge, termed Universal Differential Equations [18, 19], are an attractive modeling platform for these problems because of their flexibility in encoding structural priors.
@@ I-B2 Data-driven Modeling of DAEs
While the treatment of domain-specific problems differs, all rely on multi-step integration techniques separating differential and algebraic state updates. Recognizing this common structure, we recast these methods under operator splitting, where integration is decomposed into sequential sub-tasks. We propose a new class of neural timesteppers for data-driven modeling of DAEs with partially unknown dynamics.
@@ II-B Neural Differential Equations
where the superscript contained in parentheses denotes indexing by time. This integration is performed with a standard numerical ODE solver, which we denote as $x^{(t+\Delta t)}=\operatorname{ODESolve}\left(f,x^{(t)};\theta\right)$. This is achieved through either (i) backpropagating residuals through the elementary functions of the explicit time-stepping integrator (reverse-mode autodiff) [27, 4], (ii) an adjoint sensitivity definition and solution [7], or (iii) forward-mode autodiff [28].
II-C Operator Splitting Methods
In operator splitting methods (OS), also called fractional step methods, the evolution of a system is split into simpler, independent sub-tasks. The underlying ODE is then solved explicitly as the sequential solution of these sub-tasks, albeit at a decreased level of accuracy (i.e., splitting error [29]). There are several well-established OS schemes. The simplest one, Lie–Trotter operator splitting [5], assumes that the vector field $f$ can be decomposed into two operators:
$\frac{dx}{dt}=A(x)+B(x),$ (4)
for the interval $t\in[t,t+\Delta t]$. Where $A$ and $B$ are assumed to be locally Lipschitz continuous vector fields, ensuring the existence and uniqueness of solutions for each sub-step and the well-posedness of the splitting scheme. The Lie–Trotter splitting scheme defines the sub-tasks as:
$\hat{x}^{(t+\Delta t)}$ $=\operatorname{ODESolve}\left(A(x),x^{(t)}\right),$ (5a)
$x^{(t+\Delta t)}$ $=\operatorname{ODESolve}\left(B(x),\hat{x}^{(t+\Delta t)}\right).$ (5b)
The first sub-task solves the ODE defined by the operator $A(x)$ on the time interval $[t,t+\Delta t]$. The second sub-task solves the ODE defined by the operator $B(x)$ on the same time interval, but with initial conditions given by the result of the first sub-task. Many variations of this scheme exist for different applications and orders of accuracy. For example, Strang splitting [30] is a second-order scheme that uses half-steps (i.e., steps of $\frac{\Delta t}{2}$) in the above sequence.
@@ III PROBLEM FORMULATION
A fundamental assumption of a NODE (2) is that a direct relationship exists between each system state and its temporal derivative. This is a reasonable assumption for many cases, but poses challenges for partially-observed systems [31] and application to DAEs. Index reduction is a reliable technique for recasting a DAE into a set of ODEs for standard numerical integration [32], but this is incompatible with the problem when the DAE governing equations are unknown or incomplete, leading to an unknown system of ODEs and required state augmentations.
@@
$\underset{\theta_{f}}{\text{min}}$ $\int_{t_{0}}^{t_{N}}\Big{(}||x-\hat{x}||_{2}^{2}+||y-\hat{y}||_{2}^{2}\Big{)}$ (6a)
s.t $\frac{dx}{dt}=f(x(t),y(t),u(t);\theta_{f}),$ (6b)
$0=g(x(t),y(t),u(t)),$ (6c)
@@
Where $\hat{x},\hat{y}$ represent reference state variables generated by the unknown DAE system (1). The objective of the problem (6a) is to train the parameters $\theta_{f}$ such that the reference state trajectories in the time interval $[{t_{0}},{t_{N}}]$ are closely approximated by the NDAE model given via Eq. (6b) and Eq. (6d), respectively.
@@
Under the above assumptions, there exists a unique, continuously differentiable solution $(x^{(t)},y^{(t)})$ defined on an interval $[t_{0},t_{0}+\Delta t)$, for some $\Delta t>0$, such that the DAE system is satisfied and $(x^{(t_{0})},y^{(t_{0})})=(x_{0},y_{0})$. Furthermore, the algebraic constraint can be locally solved for $y$ as a function of $x$ and $u$, thereby reducing DAEs into a system of ODEs $\frac{\partial x}{\partial t}=f(x,h(x,u),u)$ to which the classical Picard-Lindelöf Theorem [34] applies.
Remark. Theorem III.1 follows directly from the Implicit Function Theorem under Assumption 1, which guarantees that the DAE can be reduced to an index-1 form locally. We include this result for completeness and to formalize the assumptions under which the algebraic surrogate $h$ is valid.
@@ IV METHOD
IV-A Operator Splitting for Differential Algebraic Equations
Now consider the DAE system of Eq. (1). Similar to the approach of a Lie-Trotter splitting scheme, we split the evolution of a system’s differential and algebraic states into sequential sub-tasks. The first sub-task updates the algebraic states using a trainable deep neural network (DNN). The second sub-task takes as input the updated algebraic states and performs a single-step integration. These two sequential sub-tasks define one complete forward pass; all system states are updated exactly once. More formally, consider a differentiable ODE solver, denoted as $\operatorname{ODESolve}$. Using operator splitting to solve the semi-explicit DAE in Eq. (1), we define the following neural time-stepper:
$y^{(t+\Delta t)}$ $=h\left(x^{(t)},y^{(t)},u^{(t)};\theta_{h}\right),$ (7a)
@@ IV-A Operator Splitting for Differential Algebraic Equations
where $h:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{y}}\times\mathbb{R}^{n_{u}}\rightarrow\mathbb{R}^{n_{y}}$ is a neural surrogate for the algebraic update. In classical DAE theory, such a mapping exists locally under the Implicit Function Theorem, provided that $\frac{\partial g}{\partial y}$ is nonsingular (see Assumption 1). This guarantees that, for fixed $x$ and $u$, the algebraic variable $y$ can be implicitly expressed as a function $h(x,u;\theta_{h})$ satisfying $g(x,h(x,u;\theta_{h}),u)=0$. In our framework, we approximate this mapping $h$ using a neural network and train its parameters $\theta_{h}$ to minimize the residual $\|g(x,y,u)\|$. This surrogate enables efficient algebraic updates in the first step of the operator splitting scheme. In contrast to PINNs [35], which minimize residuals for all equations simultaneously, our method separates the algebraic and differential components, handling the latter through the $\operatorname{ODESolve}$ step. This modular structure improves interpretability and computational flexibility when modeling partially known DAEs.
We denote this complete set of operations as $\operatorname{DAESolve}$, with one complete time step captured as:
@@
where $\Theta=\{\theta_{h},\theta_{f}\}$ is the set of all model parameters, and the temporal integration bounds are assumed to be $[t,t+\Delta t]$.
Figure 1 depicts the handling of differential and algebraic states through the proposed OS-based NDAE integration scheme.
@@
In contrast, we adopt an operator splitting approach with learned algebraic surrogates and explicit time integration for DAE systems. Note that in the proposed operator splitting scheme (7), updates to the algebraic and differential states are performed sequentially: the algebraic state $y(t+\Delta t)$ is first updated via the algebra surrogate $h$, and the resulting value is then used to integrate the differential state $x(t+\Delta t)$ through the $\operatorname{ODESolve}$ step. This structure is reminiscent of Neural Controlled Differential Equations [36], where external signals influence the dynamics via sequential updates. This ordering is non-commutative and plays a critical role in preserving the structure of the underlying DAE. The rationale stems from the formulation of semi-explicit DAEs (1), where the algebraic variable $y$ is not governed by a differential equation but is implicitly constrained to satisfy $g(x,y,u)=0$ at all times. As such, the algebraic state must be updated first to ensure that the subsequent evaluation of the differential dynamics $f(x,y,u)$ begins from a consistent point on or near the constraint manifold. Adopting this algebraic-first update reduces the constraint violations and aligns with standard practices in DAE solvers, which typically resolve algebraic updates before time-stepping [33].
Alternatives to the proposed scheme exist; most recently [37] proposed a reverse-order update strategy in the form of a predictor-corrector scheme, where the differential state is first propagated using an unconstrained neural ODE model, and the result is then projected onto the constraint manifold to restore algebraic consistency. Alternatively, simultaneous updates as recently proposed by [38], i.e., solving for $(x,y)$ jointly at each time step, reduce splitting error and provide tighter constraint satisfaction. However, this approach is more computationally demanding as it requires solving a constrained nonlinear optimization problem.
IV-B Loss Function and Optimization Algorithm
@@
We optimize model parameters $\Theta$ with an off-the-shelf stochastic gradient-based optimizer [39] to solve $\Theta^{*}=\underset{\Theta}{\text{argmin}}~{}\mathcal{L}(\Theta)$. For this work, we choose minimal architectures for ease of implementation and exposition, as outlined in Alg. 1. We use a first-order explicit timestepper for the ODE integration and use a DNN as a neural surrogate of an algebra solver. However, the proposed OS-based NDAE scheme is more generic than that and allows the combination of higher-order ODE solvers with iterative algebraic solvers.
@@ 23             $\Theta\longleftarrow\Theta-\eta\nabla\mathcal{L}(\Theta)$
We acknowledge that the convergence of the residual loss $\mathcal{L}(\theta)$ is not guaranteed in finite epochs, especially under noise or model mismatch. However, empirical studies suggest that with overparameterized networks and sufficient data, residual minimization can converge [40], though adding data may not always improve performance due to inconsistency or overfitting. The neural surrogates $h$ are trained to approximate solutions to the algebraic constraint in a region where the DAE remains index-1, as ensured by Assumption 1. While Theorem III.1 assumes consistent initialization, in practice, we enforce approximate consistency by penalizing initial constraint violations via a residual loss. This encourages alignment with the constraint manifold over time, even under noise or partial observability. Nevertheless, poorly trained surrogates may violate this regularity. Structural priors, as used in our case studies, or post-projection methods [37], can help ensure stable updates. These limitations motivate future work on robust residual weighting, adaptive stopping criteria, and formal convergence guarantees.
@@ V-A1 Problem Description
A manifold is a physical realization of a conservation relationship that manifests as an algebraic constraint in a DAE. Figure 2 depicts a pair of tanks fed by a manifold. A pseudo-infinite reservoir supplies the pump. At the manifold, the inflow is exactly equal to the sum of the outflows; that is, $y_{\text{in}}=y_{1}+y_{2}$, where $y$ represents volumetric flow rates and the subscripts $1$ and $2$ denote the two exit streams of the manifold. The two tanks have different area-height profiles but share a common datum such that the column heights are identical. As a fluid is pumped into the two tanks, the flow rates in the two streams vary to ensure (i) the conservation relationship of the manifold and (ii) identical pressure head of the tanks, i.e, $x_{1}=x_{2}$, where $x$ is the height of the fluid. The full DAE that describes the evolution of this system is given as:
$\begin{split}\frac{dx_{1}}{dt}=\frac{y_{1}}{\phi_{1}(x_{1})},\quad\frac{dx_{2}}{dt}=\frac{y_{2}}{\phi_{2}(x_{2})},\end{split}$ (10)
$0=y_{\text{in}}-y_{1}-y_{2},\quad 0=x_{1}-x_{2},$
where $\phi_{1}$ and $\phi_{2}$ denote the area-height profiles for the two tanks. For this study, the area-height profiles are set to $\phi_{1}(x)=3.0$ and $\phi_{2}(x)=\sqrt{x}+0.1$ respectively. The goal of the learning task is to recover $\phi_{2}(x)$ given observations of tank heights $x$ and pump volumetric flow rates $y$. Note that if the task were to recover both $\phi_{1}(x)$ and $\phi_{2}(x)$ simultaneously, the problem would be ill-posed in the sense that there does not exist a unique set of tank profiles that recovers the behavior seen in observations.
@@
Figure 2: Tank-manifold-pump schematic.